OpenStudy (nuccioreggie):

A spinner is divided into many sections of equal size. Some sections are red, some are blue, and the remaining are green. The probability of the arrow landing on a section colored red is 11 over 20. The probability of the arrow landing on a section colored blue is 6 over 20. What is the probability of the arrow landing on a green-colored section? 6 over 20 3 over 20 9 over 20 17 over 20

Mehek (mehek14):

\(\dfrac{11}{20}+\dfrac{6}{20}=\dfrac{17}{20}\)<- probability of not getting green

OpenStudy (nuccioreggie):

It B?

Mehek (mehek14):

\(\dfrac{20}{20}-\dfrac{17}{20}= probability~of~getting~green\)

Mehek (mehek14):

yes B

OpenStudy (nuccioreggie):

Ok thank you can you help me with like three or 4 more please ill tag you

Mehek (mehek14):

I gtg somewhere I'll come on in like 5-10 mins

OpenStudy (nuccioreggie):

Ok jus message me when you are ready k?